0 551 [10] b) Calcule is probabilidad de que x > 0.5, suponiendo que o = 2
Answer #1

solution t1 = ut (1+0) ocasa E691 = *5(1+0) du = (1+0) ( - 140 6044= (1+0)]26 du = (40) (x2 )! E(x)=148 94 8=0.2 VIXI = E1 2thank you